waht is the product of {x/4+2y/3}{x/4-2y/3}
Answer by sdmmadam@yahoo.com(530)

waht is the product of {x/4+2y/3}{x/4-2y/3}
On studying the problem, two quantities x/4 and 2y/3 occur in both the brackets.
While the first bracket is the sum of the two quantities,
the second bracket is the difference.
This should remind us of the product (a+b)(a-b) which is a^2 - b^2 (either by direct multiplication or by formula)
Thus using (a+b)(a-b)= a^2 - b^2
we have {x/4+2y/3}{x/4-2y/3}
= (x/4)^2 - (2y/3)^2 [here a= (x/4) and b = (2y/3)]
= x^2/16 - 4y^2/9